OpenStudy (anonymous):

how to find tangent line of y=cosx-sinx at (pi, -1)

OpenStudy (anonymous):

since the point is on the curve, find dy/dx using (pi,-1) to find the equation of the line

OpenStudy (anonymous):

Take the derivative y'=-sinx-cosx, evaluate at (pi,-1) y' at (pi,-1)= 0 - (-1) = 1 = slope m of the tangent line Point-slope y-(-1)=(1)(x-pi) Therefore y=x-pi-1

OpenStudy (anonymous):

lim {[(cosx+h)-sin(x+h)]-[cosx-sinx]}/h as h->0